### v1

# Kleene Algebrae, Kleene Modules, and Morita Equivalence

Kleene Algebrae, Kleene Modules, and Morita Equivalence
Luke Serafin (lss255)
(October 28, 2025)
1 Introduction
Modules and the notion of Morita equivalence are foundational to the classical study of rings. These concepts extend naturally to semirings and then specialize to Kleene algebrae, and my goal is to investigate Kleene modules and Morita equivalence of Kleene algebrae in the hope that some of the power seen in the context of rings may be found in this new context as well. I undertook the definition and study of Kleene modules on my own, but was aware that others had done so and came across the notion in a form—which is exactly what I settled on—in [1]. I was not until recently aware of the fact that Morita equivalence has been studied in the context of semirings, but it has. See for example [2].
2 Kleene modules
Definition 1.
Let $K$ be a Kleene algebra. A left Kleene module over $K$ is a left idempotent semiring module $M$ over the semiring reduct of $K$ which additionally satisfies that $am\leq m$ implies $a^{*}m\leq m$ for $m\in M$ and $a\in K$. A right Kleene module is similarly a right module satisfying that $ma\leq m$ implies $ma^{*}\leq m$. A Kleene $(K,L)$-bimodule is a structure which is simultaneously a left Kleene $K$-module and a right Kleene $L$-module, such that if $m\in M$, $a\in K$, and $b\in L$, then $(am)b=a(mb)$.
Kleene modules are a quasivariety, and so by universal algebra there are free Kleene $K$-modules over arbitrary sets of generators for $K$ a fixed Kleene algebra.
Examples
Algebrae
Any Kleene algebra $K$ is a $(K,K)$-bimodule over itself, and a $(2,2)$-bimodule over the Kleene algebra $2$ with two elements. Moreover it is an $(A,A)$-bimodule over any subalgebra $A$ of $K$. In particular, if $K$ is a Kleene algebra with tests (a KAT), and $B$ is its distinguished Boolean subalgebra, then $B$ is a Kleene subalgebra of $K$ and $K$ is a $(B,B)$-bimodule.
Ideals
Let $I$ be a left ideal of a Kleene algebra $K$. Then $I$ is a left Kleene module with action given by multiplication. To see this, note that $I$ is clearly a left idempotent semiring module under the action of multiplication, and that for $m\in I$, if $am\leq m$ then by the *-axioms $a^{*}m\leq m$. Similary, if $J$ is a right ideal of $K$ then $J$ is a right Kleene module, and if $J$ is a two-sided ideal then it is a $(K,K)$-Kleene bimodule.
Quotients
Let $\equiv$ be a congruence of a Kleene algebra $K$, and set $Q=K/{\equiv}$. Then $Q$ is a Kleene $K$-bimodule with action given by multiplication. This follows from the fact that $\equiv$ is a congruence, and provides an example of a Kleene algebra with a natural Kleene bimodule structure.
Function modules
Let $K$ be a Kleene algebra, $B$ be a set, and consider the set $K^{B}$ of functions from $B$ to $K$ endowed with pointwise addition and zero as the constant-zero function. Define left and right actions of $K$ on $K^{B}$ by $(am)(b)=am(b)$ and $(ma)(b)=m(b)a$ for $a\in K$, $m\in K^{B}$. It is immediate that these operations give a $(K,K)$-bimodule structure on $K^{B}$. Moreover, if $B$ is finite, then $K^{B}$ has the usual finite basis given by characteristic functions for singletons in $B$, and is the free $(K,K)$-bimodule with generators $B$. Forgetting either the right or the left action, it is also the free left and the free right $K$-module with generators $B$. One can also restrict from $K^{B}$ to functions supported on a set in some lattice-ideal for $\operatorname{\mathcal{P}}(B)$. It is easy to see that the finitely-supported elements of $K^{B}$ form the free $K$-module with generators $B$.
Homomorphism modules
Fix a Kleene algebra $K$ and let $M$, NN be left $K$-modules. The set $\operatorname{Hom}(M,N)$ becomes a left $K$-module when endowed with pointwise operations. Similarly for $M$, NN right $K$-modules, or $(K,S)$-bimodules where $S$ is another Kleene algebra.
Endomorphism modules
Let $x\in K$, $\alpha\in\mathop{\text{End}}(M)$. If $x\alpha\leq\alpha$, then for every $k\in K$, $x\alpha(k)\leq\alpha(k)$, so $x^{*}\alpha(k)\leq\alpha(k)$ for every $k\in K$ and hence $x^{*}\alpha\leq\alpha$. Consequently $\mathop{\text{End}}(K)$ is a left Kleene module under multiplication by elements of $K$. Similarly it is a right Kleene module under multiplication by elements of $K$, and it is easy to check that it is a $(K,K)$-bimodule. Matrices are the special case of this where $M=K^{n}$.
Propositional Dynamic Logic
Fix sets $P$ of action atoms and $T$ of test atoms, respectively. Let $K$ be the free Kleene algebra generated by $P$, and $B$ the free Boolean algebra generated by $T$. The Lindenbaum-Tarski algebra $A$ of propositional dynamic logic with action atoms $P$ and test atoms $T$ can be viewed as a $(B,B)$-bimodule via the subalgebra bimodule structure, but also has the structure of a left $K$-module via the action $pa=[\langle p\rangle\phi]$ where $p\in K$, $a\in A$, $a=[\phi]$, and brackets denote the equivalence class of a formula in the Lindenbaum-Tarski algebra. This follows immediately from the axioms for propositional dynamic logic.
The dual of a Kleene module
Definition 2.
Fix a Kleene algebra $K$, and let $M$ be a left (respectively, right) Kleene module over $K$. The dual $M^{\circ}$ of $M$ is the set of Kleene $K$-module homomorphisms from $M$ to $K$, with pointwise addition and 0 the constant 0-function.
The proof that the dual of a Kleene module is a Kleene module is immediate.
Proposition 1.
For $M$ a left (respectively, right) Kleene module over $K$, its dual $M^{\circ}$ is a right (respectively, left) Kleene module over $K$. Moreover the dual of a Kleene $(K,S)$-bimodule is a Kleene $(S,K)$-bimodule.
Proof.
The dual of $M$ is isomorphic to the homomorphism module $\operatorname{Hom}(M,K)$. ∎
Tensor products of Kleene modules
Definition 3.
Fix Kleene algebrae $A$, $B$, $C$. A tensor product is a functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$ such that for each $N\in{{}_{B}\text{Mod}_{C}}$, the functors $\lambda M.\,F(M,N)$ and $\lambda P.\,\operatorname{Hom}(N,P)$ are adjoint.
Note that it follows from the universal property for adjoints that tensor products, if they exist, are unique up to isomorphism. One can also define tensor products producing left and right Kleene modules, with types ${{}_{A}\text{Mod}_{B}}\times\text{Mod}_{B}\rightarrow\text{Mod}_{A}$ and ${{}_{A}\text{Mod}}\times{{}_{A}\text{Mod}_{B}}\rightarrow{{}_{A}\text{Mod}}$, in the obvious way. We omit the definition since it won’t be used in the sequel.
Defining tensor products of Kleene modules doesn’t guarantee they exist; this is established next.
Proposition 2.
For any Kleene algebrae $A$, $B$, $C$ there is a tensor product functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$.
Proof.
We proceed by direct construction. Fix $M\in{{}_{A}\text{Mod}_{B}}$ and $N\in{{}_{B}\text{Mod}_{C}}$. Let $\Xi$ be the Kleene module freely generated by $M\times N$, and let $\Phi$ be the congruence generated by the relations
$(m+m^{\prime},n)\Phi(m,n)+(m^{\prime},n),(m,n+n^{\prime})\Phi(m,n)+(m,n^{\prime}),(mb,n)\Phi(m,bn)$
for $m,m^{\prime}\in M$, $n,n^{\prime}\in N$, and $b\in B$. Define $M\otimes_{B}N=\Xi/\Phi$.
It suffices to establish that the functors $M\mapsto M\otimes_{B}N$ and $P\mapsto\,\operatorname{Hom}(N,P)$ are adjoint. This means that there is a natural isomorphism $\operatorname{Hom}(M\otimes_{B}N,P)\cong\operatorname{Hom}(M,\operatorname{Hom}(N,P))$ for $M\in{{}_{A}\text{Mod}_{B}}$ and $N,P\in{{}_{B}\text{Mod}_{C}}$. For $\varphi\in\operatorname{Hom}(M\otimes_{B}N,P)$ define $\overline{\varphi}$ by $\overline{\varphi}(m)(n)=\varphi(m\otimes n)$. This defines a homomorphism $M\rightarrow\operatorname{Hom}(N,P)$ by freeness of $\Xi$ and the construction of the congruence $\Phi$. Let $M^{\prime}\in{{}_{A}\text{Mod}_{B}}$, $P^{\prime}\in{{}_{B}\text{Mod}_{C}}$, $\alpha\mathrel{:}M^{\prime}\rightarrow M$, and $\beta\mathrel{:}P\rightarrow P^{\prime}$. To see that the map $\varphi\mapsto\overline{\varphi}$ is a natural transformation, we must verify that the following diagram commutes, where the horizontal arrows represent application of the transformation $\varphi\mapsto\overline{\varphi}$.
$\operatorname{Hom}(M\otimes_{B}N,P)$$\operatorname{Hom}(M,\operatorname{Hom}(N,P))$$\operatorname{Hom}(M^{\prime}\otimes_{B}N,P^{\prime})$$\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$$\scriptstyle{\operatorname{Hom}(\alpha\otimes\text{id},\alpha)}$$\scriptstyle{\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))}$
Let $\varphi\in\operatorname{Hom}(M\otimes_{B}N,P)$. Then for $m^{\prime}\in M^{\prime}$ and $n\in N$,
$\overline{\operatorname{Hom}(\alpha\otimes\text{id},\beta)(\varphi)}(m^{\prime})(n)$ $=(\beta\circ\varphi\circ(\alpha\otimes\text{id}))(m^{\prime}\otimes n)$
$=\beta(\varphi((\alpha\otimes\text{id})(m^{\prime}\otimes n)))$
$=\beta(\varphi(\alpha(m^{\prime})\otimes n))$
and
$\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))(\overline{\varphi})(m^{\prime})(n)$ $=\operatorname{Hom}(N,\beta)(\overline{\varphi}(\alpha(m^{\prime}))(n))$
$=\operatorname{Hom}(N,\beta)(\varphi(\alpha(m^{\prime})\otimes n))$
$=\beta(\varphi(\alpha(m^{\prime})\otimes n))$
Consequently $\varphi\mapsto\overline{\varphi}$ is a natural transformation.
It remains to find an inverse to the natural transformation $\varphi\mapsto\overline{\varphi}$. The natural candidate is given for $M\in{{}_{A}\text{Mod}_{B}}$, $P\in{{}_{B}\text{Mod}_{C}}$, $\psi\in\operatorname{Hom}(M,\operatorname{Hom}(N,P))$, $m\in M$, and $n\in N$ by $\hat{\psi}(m\otimes n)=\psi(m)(n)$. It is immediate that the transformation $\psi\mapsto\hat{\psi}$ is inverse to the transformation $\phi\mapsto\overline{\phi}$. It remains to verify that $\psi\mapsto\hat{\psi}$ is a natural transformation. For this we need to check that the following diagram commutes, where $M,M^{\prime}\in{{}_{A}\text{Mod}_{B}}$, $P,P^{\prime}\in{{}_{B}\text{Mod}_{C}}$, $\alpha\mathrel{:}M\rightarrow M^{\prime}$, $\beta\mathrel{:}P^{\prime}\rightarrow P$, and the horizontal arrows are given by application of the transformation $\psi\mapsto\hat{\psi}$.
$\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$$\operatorname{Hom}(M^{\prime}\otimes_{B}N,P^{\prime})$$\operatorname{Hom}(M,\operatorname{Hom}(N,P))$$\operatorname{Hom}(M\otimes_{B}N,P)$$\scriptstyle{\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))}$$\scriptstyle{\operatorname{Hom}(\alpha\otimes\text{id},\beta)}$
Let $\psi\in\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$. Then for $m\otimes n\in M\otimes_{B}N$,
$\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))\hat{\phantom{|}}(\psi)(m\otimes n)$ $=\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))(\psi)(m)(n)$
$=(\operatorname{Hom}(N,\beta)\circ\psi\circ\alpha)(m)(n)$
$=\operatorname{Hom}(N,\beta)(\psi(\alpha(m)))(n)$
$=(\beta\circ\psi(\alpha(m)))(n)$
$=\beta(\psi(\alpha(m))(n))$
One more computation now verifies that this diagram commutes:
$\operatorname{Hom}(\alpha\otimes\text{id},\beta)(\hat{\psi})(m\otimes n)$ $=(\beta\circ\hat{\psi}\circ(\alpha\otimes\text{id}))(m\otimes n)$
$=\beta(\hat{\psi}(\alpha(m)\otimes n))$
$=\beta(\psi(\alpha(m))(n)).$
Therefore the transformation $\psi\mapsto\hat{\psi}$ is inverse to the natural transformation $\varphi\mapsto\overline{\varphi}$, so these transformations are natural isomorphisms. ∎
2.1 The Morita Category
Morita equivalence is of fundamental importance to ring theory, and based on the suspicion that it may be useful in the study of Kleene algebrae as well we develop it here. Morita equivalence for rings is obtained by replacing the morphisms of the category of rings by bimodules, and considering the induced notion of isomorphism. More specifically, the morphisms from a ring $R$ to a ring $S$ are the $(R,S)$-bimodules, and composition of bimodules is given by tensor product. For more on Morita equivalence of rings see [3].
The case of Morita equivalence for Kleene algebrae is a straightforward adaptation of that for rings.
Definition 4.
The Morita category of Kleene algebrae has objects Kleene algebrae, and for $A$ and $B$ Kleene algebrae the morphisms between them are all Kleene $(A,B)$-bimodules. The identity at an algebra $K$ is ${{}_{K}}K_{K}$, and composition of morphisms is given by tensor product.
To verify that this is indeed a category, we must check that tensor product is associative and has the bimodules ${{}_{K}}K_{K}$ as identities.
Proposition 3.
Let $A$, $B$, $C$, $D$ be Kleene algebrae, and ${{}_{A}}M_{B}$, ${{}_{B}}N_{C}$, ${{}_{C}}P_{D}$ be Kleene bimodules. Then
$({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D}\cong{{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D}),$
and also ${{}_{A}}A_{A}\otimes{{}_{A}}M_{B}\cong{{}_{A}}M_{B}$ and ${{}_{A}}M_{B}\otimes{{}_{B}}B_{B}\cong{{}_{A}}M_{B}$.
Proof.
By the defining property of tensor products, for any Kleene module ${{}_{D}}Q_{E}$ we natural isomorphisms
$\operatorname{Hom}(({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D},{{}_{D}}Q_{E})\cong\operatorname{Hom}(M,\operatorname{Hom}(N,\operatorname{Hom}(P,Q)))$
and
$\operatorname{Hom}({{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D}),{{}_{D}}Q_{E})\cong\operatorname{Hom}(M,\operatorname{Hom}(N,\operatorname{Hom}(P,Q))).$
Hence $({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D}\cong{{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D})$. Also for any $a\in A$ and $m\in M$, $a\otimes m=a(1\otimes m)$, so the map $m\mapsto 1\otimes m$ is an isomorphism from $M$ to $A\otimes M$. ∎
As we have defined it, the Morita “category” isn’t actually a category, because the associativity and identity laws only hold up to isomorphism. In practice this won’t matter, since isomorphic Kleene algebrae and modules are regarded as “the same.” If one insists on working with a bona fide category, one may restrict to a skeleton of the Morita category, which includes precisely one element of each isomorphism class of objects and morphisms from the Morita category. Also, we said that the Morita category extends the usual category of Kleene algebrae, and must check that each homomorphism between Kleene algebrae gives rise to a Kleene module in a manner which sends a composition of homomorphisms to the tensor product of the corresponding modules.
Definition 5.
Let $A$, $B$ be Kleene algebrae and $h\mathrel{:}A\rightarrow B$ be a homomorphism. The homomorphism module $E_{h}$ is a Kleene $(A,B)$-bimodule whose elements and addition are given by the additive reduct of $B$, with the following multiplication actions on $b\in B$: For $a\in A$, $ab=h(a)b$. For $^{\prime}\in B, bb$$b^{\prime}=bb^{\prime}$ (just perform the multiplication in $B$). It is immediate that this givs $E_{h}$ the structure of a Kleene bimodule.
Proposition 4.
Let $A$, $B$, $C$ be Kleene algebrae and $f\mathrel{:}A\rightarrow B$ and $g\mathrel{:}B\rightarrow C$ be homomorphisms. Then $E_{g\circ f}\cong E_{f}\otimes E_{g}$.
Proof.
We shall show that the map $\phi\mathrel{:}E_{g\circ f}\rightarrow E_{f}\otimes E_{g}$ given by $\phi(c)=1\otimes c$ is an isomorphism. To see that this is a homomorphism, additivity and the commutativity with the right action of $C$ are immediate. For the left action of $A$, note that for $a\in A$,
$\phi(ac)$ $=1\otimes(g\circ f)(a)c$
$=f(a)\otimes c$
$=a(1\otimes c)=a\phi(c).$
Now consider $b\otimes c\in E_{g}\otimes E_{f}$. Define $\psi\mathrel{:}E_{f}\otimes E_{g}\rightarrow E_{g\circ f}$ by $\psi(b\otimes c)=bc$, and note that this is well-defined because $b\otimes c=1\otimes bc$ for $b\in B$, $c\in C$. It is clearly additive and commutes with the right action of $C$, so we demonstrate that it commutes with the left action of $A$. Let $a\in A$ and note that
$\psi(a(b\otimes c))$ $=\psi(f(a)(b\otimes c))$
$=\psi(f(a)\otimes bc)$
$=f(a)bc$
$=a\psi(b\otimes c)$
It remains to compute
$\phi(\psi(b\otimes c))=\phi(bc)=1\otimes bc=b\otimes c$
and
$\psi(\phi(c))=\psi(1\otimes c)=1c=c$
to see that $\phi$ and $\psi$ are inverse bimodule homomorphisms, hence isomorphisms. Therefore $E_{g\circ f}\cong E_{f}\otimes E_{g}$. ∎
Another way of expressing Morita equivalence for rings is as equivalence of the corresponding module categories. This carries over to Kleene algebrae, as we now argue. If $K$ is Morita-equivalent with $S$ via bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$, then for any left $K$-module ${{}_{K}}P$,
$M\otimes(N\otimes P)\cong(M\otimes N)\otimes P\cong K\otimes P\cong P,$
and for any left $S$-module ${{}_{S}}Q$,
$N\otimes(M\otimes Q)\cong(N\otimes M)\otimes Q\cong S\otimes Q\cong Q,$
so the functors $P\mapsto N\otimes P$ from $\text{Mod}_{K}$ to $\text{Mod}_{S}$, and $Q\mapsto N\otimes Q$ from $\text{Mod}_{S}$ to $\text{Mod}_{K}$, constitute an equivalence of categories. The right-module categories $_{K}\text{Mod}$} and $_{S}\text{Mod}$} are equivalent by an exactly symmetric argument, and combining these arguments together gives that the bimodule categories $_{K}\text{Mod}_{S}$} and $_{S}\text{Mod}_{K}$} are equivalent.
As in the classical case, a Kleene algebra is Morita-equivalent with each of its finite-dimensional matrix rings. This can be viewed as explaining why the behaviour of finite automata over a Kleene algebrae, which can be defined by finite-dimensional square matrices, is fully expressible in terms of the base Kleene algebra.
Proposition 5.
Let $K$ be a Kleene algebra, $n$ a positive integer, and $M$ be the Kleene algebra of $n\times n$ matrices over $K$. Then $K$ is Morita-equivalent with $M$.
Proof.
The $(M,K)$-bimodule $K^{n}$ and its dual $K^{n\circ}$ will witness this equivalence. First consider $K^{n}\otimes K^{n\circ}$. For $1\leq i\leq n$ let $e_{i}$ be the tuple which is zero everywhere except position $i$, where it is $1$. Note that the set $\{e_{i}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n}$. Let $e_{i}^{\circ}$ be the dual of this module element, that is, $e_{i}^{\circ}(e_{i})=1$ and $e_{i}^{\circ}(e_{j})=0$ for $j\neq i$. It is easy to show that the set $\{e_{i}^{\circ}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n\circ}$. For $1\leq i,j\leq n$, let $E_{ij}\in M$ be the matrix which is zero everywhere except at position $(i,j)$, where it is $1$. Note that these freely generate $M$, and that the set $\{e_{i}\otimes e_{j}^{\circ}\mathrel{:}1\leq i,j\leq n\}$ freely generates $K^{n}\otimes K^{n\circ}$, because only elements of $K$ can be transported across the tensor product. Define a map $\phi\mathrel{:}K^{n}\otimes K^{n\circ}$ by $\phi(e_{i}\otimes e_{j}^{\circ})=E_{ij}$; we show that $\phi$ is an isomorphism. Its inverse, naturally, will be given by $\psi(E_{ij})=e_{i}\otimes e_{j}^{\circ}$. By free generation $\phi$ and $\psi$ are well-defined homomorphisms, and it is immediate that they are inverses. Consequently $K^{n}\otimes K^{n\circ}\cong M$, which is half of a Morita equivalence.
For the other half we look at $K^{n\circ}\otimes K^{n}$, and our goal is to show this is isomorphic to $K$. For $e_{i}^{\circ}\otimes e_{j}\in K^{n\circ}\otimes K^{n}$, define $\alpha(e_{i}^{\circ}\otimes e_{j})$ to be 0 if $i\neq j$ and $1$ if $i=j$. By free generation this determines a homomorphism $\alpha\mathrel{:}K^{n\circ}\otimes K^{n}\rightarrow K$. Define also $\beta\mathrel{:}K\rightarrow K^{n\circ}\otimes K^{n}$ by $\beta(a)=a(e_{1}^{\circ}\otimes e_{1})$ for $a\in A$, which is well-defined by free generation. To complete the proof we show that $\alpha$ and $\beta$ are inverses. Certainly for $a\in A$ we have that
$\alpha(\beta(a))=\alpha(a(e_{1}^{\circ}\otimes e_{1}))=a\alpha(e_{1}^{\circ}\otimes e_{1})=a1=a.$
To see that $\beta\circ\alpha=\text{id}$ it suffices to show that for any $i\leq n$, $e_{i}^{\circ}\otimes e_{i}=e_{1}^{\circ}\otimes e_{1}$. For this we show that
$e_{i}^{\circ}\otimes e_{i}=\left(\sum_{i=1}^{n}e_{i}^{\circ}\right)\otimes\left(\sum_{i=1}^{n}e_{i}\right).$
Before writing this out in generality, an example might be instructive. We compute in the case $n=3$ as follows:
$\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$ $=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
To compute this in generality, let $\overline{1}$ be the $n\times n$ matrix with every entry $1$, and $\overline{e}_{i}$ be the result of replacing the $i$th row of $\overline{1}$ with the entries of $e_{i}$. It is straightforward to compute the following:
  • $e_{i}^{\circ}\overline{e}_{i}=e_{i}^{\circ}$.
  • $\overline{e}_{i}e_{i}=\sum_{i=1}^{n}e_{i}$.
  • $\overline{1}\sum_{i=1}^{n}e_{i}=\sum_{i=1}^{n}e_{i}$.
  • $e_{i}^{\circ}\overline{1}=\sum_{i=1}^{n}e_{i}^{\circ}$.
Then we have
$e_{i}^{\circ}\otimes e_{i}$ $=e_{i}^{\circ}\overline{e}_{i}\otimes e_{i}$
$=e_{i}^{\circ}\otimes\overline{e}_{i}e_{i}$
$=e_{i}^{\circ}\otimes\sum_{i=1}^{n}e_{i}$
$=e_{i}^{\circ}\otimes\overline{1}\sum_{i=1}^{n}e_{i}$
$=e_{i}^{\circ}\overline{1}\otimes\sum_{i=1}^{n}e_{i}$
$=\left(\sum_{i=1}^{n}e_{i}^{\circ}\right)\otimes\sum_{i=1}^{n}e_{i}.$
This completes the proof. ∎
There is a well-known characterization of Morita equivalence for rings which extends to semirings [3], [2]. To state it we need the following definition:
Definition 6.
Let $R$ be a semiring and $e\in R$ a multiplicative idempotent (so $e^{2}=e$). The idempotent $e$ is full if and only if $ReR=R$.
Proposition 6 ([2]).
Semirings $R$ and $S$ are Morita-equivalent if and only if for some dimension $n$ and for some full idempotent $e$ in the semiring $M$ of $n\times n$ matrices over the ring (without loss of generality) $R$, $S\cong eMe$.
Certainly Morita-equivalent Kleene algebrae are Morita equivalent as semirings, and so any Kleene algebra $S$ Morita-equivalent to the algebra $K$ is, up to isomorphism, of the form $eMe$ for some finite-dimensional matrix ring $R$ over $K$ and some full idempotent $e\in M$. It is natural to ask whether Kleene algebrae whose semiring reducts are Morita equivalent must be Morita equivalent as Kleene algebrae. That is to say, if there is a semimodule Morita isomorphism between Kleene algebrae $K$ and $S$, must there also be a Kleene algebra Morita isomorphism?
Proposition 7.
Let $K$ and $S$ be Kleene algebrae. If the semiring reducts of $K$ and $S$ are Morita equivalent, then $K$ and $S$ are Morita equivalent as Kleene algebrae.
Proof.
This appears to be hard to do directly, though maybe I just wasn’t thinking in the right direction. In any case, it is straightforward given the characterization of Morita-equivalence for semirings given above. Without loss of generality, fix a dimension $n$ such that, letting $M$ be the Kleene algebra of $n\times n$ matrices over $K$, there is a full idempotent of $M$ such that $S\cong eMe$. Then $K^{n\circ}e\otimes eK^{n}$ is a submodule of $K^{n\circ}\otimes K^{n}$, so because the idempotent $e$ is full, in fact $K^{n\circ}e\otimes eK^{n}\cong K$. Conversely,
$eK^{n}\otimes K^{n\circ}e\cong e(K^{n}\otimes K^{n\circ})e\cong eMe\cong S.$
Therefore the Kleene modules $K^{n\circ}e$ and $eK^{n}$ witness that $K$ and $M$ are Morita equivalent. ∎
Acknowledgement. Thank you to Dexter Kozen for many stimulating discussions of this material, and to Goktug for helping to finalize a project direction.