### v2

# Morita Rigidity for Kleene Algebras

Abstract
We introduce Morita equivalence to the study of Kleene algebras and modules. Classical characterizations of Morita-equivalent semirings such as having equivalent categories of modules and one semiring being a full matrix algebra over the other carry over. We also observe that Morita equivalence can be applied to extending and restricting scalars in Lindenbaum–Tarski algebras of propositional dynamic logics. But the signature result which we obtain is a form of rigidity for Kleene algebras, which states that if the semiring reducts of two Kleene algebras are Morita-equivalent, then the Morita equivalence is in fact witnessed by Kleene bimodules.
1 Introduction
Modules and the notion of Morita equivalence are foundational to the classical study of rings [15], and we assume the reader has some familiarity with the classical theory of modules though not necessarily of Morita equivalence.
The theory of modules has previously been extended to the context of semirings (for example see [5]), though we spend some time developing the theory of tensor products since these do not appear to have been previously worked out for Kleene modules, and the standard definition in [5] is inadequate for us because it makes all tensor products of modules over additively idempotent semirings (such as Kleene algebras) trivial. The alternative tensor product defined in [8] does have the properties we need, but the author had worked out what turned out to be the same definition as Katsov’s before discovering Katsov’s paper and so has included a full development of tensor products for Kleene algebras rather than simply extending Katsov’s definition for modules over semirings to the case of Kleene modules.
Additionally, a variety of instances of Kleene modules are presented, the most important being endomorphism modules of Kleene modules (including $n\times n$ matrices) and the Lindenbaum–Tarski algebra of propositional dynamic logic. Morita equivalence for rings is obtained by replacing the morphisms of the category of rings by bimodules, and considering the induced notion of isomorphism. More specifically, the morphisms from a ring $R$ to a ring $S$ are the $(R,S)$-bimodules, and composition of bimodules is given by tensor product. The author has undertaken to study Morita equivalence in the context of Kleene algebras, leading to the discovery of a number of results and directions for further study. Several standard results carry over, such as characterizations of Morita equivalence in terms of categorical equivalence of module categories, and in terms of matrix algebras defined in terms of full (multiplicative) idempotents. One surprising fact is that if the semiring reducts of two Kleene algebras $K$ and $S$ are Morita-equivalent, then in fact $K$ and $S$ are Morita-equivalent as Kleene algebras (using Kleene bimodules to witness the equivalence). This is a rigidity theorem, saying that a weaker equivalence (Morita equivalence in the category of semirings) implies a stronger equivalence (Morita equivalence in the category of Kleene algebras) within a particular domain of structures (here, Kleene algebras).
2 Semiring Preliminaries
We review here the basics of semirings and the corresponding generalization of module (sometimes called a semimodule in the literature). For more concerning semirings see for example [5] and [6]. In [5] the author develops a theory of tensor products of semirings, but these tensor products are always cancellative, which means trivial in the context of additively idempotent semirings with which we are concerned. Therefore we develop an alternative, and in the author’s opinion more natural, tensor product which is useful even in the case of additively idempotent semirings. Though this tensor product appears to be new, a form of tensor product for Kleene algebras is developed in [7], and the book [4] considers tensor products for a variety of idempotent structures related to Kleene algebras.
Definition 1.
A (unital) semiring is an algebraic structure $\langle S,+,\cdot,0,1\rangle$ such that
  1. 1.
    $\langle S,+,0\rangle$ is a commutative monoid,
  2. 2.
    $\langle S,\cdot,1\rangle$ is a monoid (not necessarily commutative),
  3. 3.
    $s\cdot 0=0\cdot s=0$ for every $s\in S$, and
  4. 4.
    multiplication distributes over addition, meaning that for any $x,y,z\in S$, $x\cdot(y+z)=(x\cdot y)+(x\cdot z)$ and $(x+y)\cdot z=(x\cdot z)+(y\cdot z)$.
We often abuse notation and denote the semiring $\langle S,+,\cdot,0,1\rangle$ by $S$, the operations being understood. Also, we generally denote multiplication $\cdot$ by juxtaposition, $xy=x\cdot y$, and regard multiplication as taking precedence over addition in expressions, so $xy+z$ is parsed as $(x\cdot y)+z$. When multiple semirings are under consideration we may add subscripts to operations, e.g. $\alpha(r+_{S}s)\cdot_{T}t$.
As is common in ring theory, we sometimes need multiplication in a semiring to be computed in the opposite order, and so define the operation on semirings which reverses the order of multiplication.
Definition 2.
For $S$ a semiring, the opposite semiring ${S}^{\mathrm{op}}$ has the same underlying set, addition, and constants as $S$, but the multiplication $\cdot_{{S}^{\mathrm{op}}}$ is defined by $s\cdot_{{S}^{\mathrm{op}}}t=t\cdot_{S}s$ for $s,t\in S$.
Note that $({S}^{\mathrm{op}})^{\mathrm{op}}=S$ for any semiring $S$. We next give the definition of a module over a semiring (also called a semimodule in the literature; e.g. [5]).
Definition 3.
Let $S$ be a semiring. A left semiring module over $S$ (or simply a left $S$-module, where it is understood that $S$ is a semiring) is a commutative monoid $\langle M,+,0\rangle$ together with a map $S\times M\rightarrow M$, denoted by juxtaposition—so $(s,m)\mapsto sm$ for $s\in S$ and $m\in M$—satisfying the following for any $s,t\in S$ and $m,n\in M$:
  1. 1.
    $s(m+n)=sm+sn$
  2. 2.
    $(s+t)m=sm+tm$
  3. 3.
    $s(tm)=(st)m$
  4. 4.
    $1m=m$
  5. 5.
    $0m=0$.
Right $S$-modules are defined symmetrically, with the operation now a map $M\times S\rightarrow M$, written on the right: $(m,s)\mapsto ms$.
For $S$ and $T$ semirings, an $(S,T)$-bimodule is a commutative monoid $\langle M,+,0\rangle$ equipped with both a left $S$-module map and a right $T$-module map, such that these two structures are compatible in the sense that for any $s\in S$, $t\in T$, and $m\in M$, $s(mt)=(sm)t$. As in the case of semirings, when we need to disambiguate in which module an operation is happening, we use subscripts, as in $s\cdot_{N}\alpha(m+_{M}n\cdot_{M}t)$ where $M$ is a right semiring module and $N$ a left semiring module over a semiring $S$. As in the literature on modules over rings, we sometimes use subscripts to the left or right of a symbol denoting a module over a semiring to indicate both which semiring acts on the module and from which side. Thus if $S$, $T$ are semirings, ${{}_{S}}M$ denotes a left module over $S$, $N_{T}$ denotes a right module over $T$, and ${{}_{S}}B_{T}$ indicates an $(S,T)$-bimodule. This notation is particularly helpful when working with direct sums and tensor products, as we shall later.
To define tensor products we shall need the notion of a congruence relation. This is a general notion from universal algebra (see for example [3]). We view modules as algebraic structures with a potentially infinite number of unary operations of scalar multiplication, one for each scalar.
Definition 4.
A congruence on an algebraic structure $\mathfrak{A}$ is an equivalence relation on the underlying set of $\mathfrak{A}$ which preserves all operations. For example, if $\mathfrak{A}$ is a semiring and $\equiv$ is a congruence on $\mathfrak{A}$, then for $a,b,a^{\prime},b^{\prime}$ elements of the underlying set of $\mathfrak{A}$, if $a\equiv a^{\prime}$ and $b\equiv b^{\prime}$, then $a+b\equiv a^{\prime}+b^{\prime}$ and $ab\equiv a^{\prime}b^{\prime}$.
It is a well-known fact from universal algebra (see [3] for a proof) that if $\equiv$ is a congruence on an algebraic structure $\mathfrak{A}$, then the set of equivalence classes of $\equiv$ carry a structure of the same type as $\mathfrak{A}$, with operations computed by taking the equivalence class of the corresponding operation in $\mathfrak{A}$ performed on representatives. This is the quotient of the algebra $\mathfrak{A}$ by the congruence $\equiv$. Thus for example if $\equiv$ is a congruence on a semiring $S$ and we use brackets to denote $\equiv$-equivalence classes, then in the quotient of $S$ by $\equiv$ we have for $s,t\in S$ that $[s]+[t]=[s+t]$ and $[s][t]=[st]$. It is easy to use the stipulation that $\equiv$ is a congruence to prove that these operations are well-defined.
As in the case of semirings, we can define the opposite of a semiring module.
Definition 5.
For $S$ a semiring and $M$ a left semiring module over $S$, the opposite module ${M}^{\mathrm{op}}$ to $M$ is a right ${S}^{\mathrm{op}}$-module with the action of ${S}^{\mathrm{op}}$ given by $m\cdot_{{M}^{\mathrm{op}}}s=s\cdot_{M}m$ for $s\in S$ and $m\in M$. The opposite of a right $S$-module and the opposite of an $(S,T)$-bimodule for $T$ also a semiring are defined similarly, with the opposite of a right $S$-module being a left ${S}^{\mathrm{op}}$-module and the opposite of an $(S,T)$-bimodule being a $({T}^{\mathrm{op}},{S}^{\mathrm{op}})$-bimodule.
Proposition 6.
The opposite of a module over a semiring $S$ is a module of the appropriate type.
Proof.
Let $S$ and $T$ be semirings; we treat just the case of an $(S,T)$-bimodule, since the other cases are straightforward adaptations of the argument. Fix an $(S,T)$-bimodule $M$ and note that the actions of ${T}^{\mathrm{op}}$ and of ${S}^{\mathrm{op}}$ on ${M}^{\mathrm{op}}$ are given respectively by
$t\cdot_{{M}^{\mathrm{op}}}m=m\cdot_{M}t\quad\text{and}\quad m\cdot_{{M}^{\mathrm{op}}}s=s\cdot_{M}m$
for $s\in S$, $t\in T$, and $m\in M$. It is immediate that these actions are additive, preserve zero, and have the property that $1_{{S}^{\mathrm{op}}}$ and $1_{{T}^{\mathrm{op}}}$ act as identities. It remains to verify that for $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, and $m\in M$, we have the following:
$t\cdot_{{M}^{\mathrm{op}}}(t^{\prime}\cdot_{{M}^{\mathrm{op}}}m)=(t\cdot_{{M}^{\mathrm{op}}}t^{\prime})\cdot_{{M}^{\mathrm{op}}}m,$ (1)
$(m\cdot_{{M}^{\mathrm{op}}}s)\cdot_{{M}^{\mathrm{op}}}s^{\prime})=m\cdot_{{M}^{\mathrm{op}}}(s\cdot_{{M}^{\mathrm{op}}}s^{\prime}),\quad\text{and}$ (2)
$(t\cdot_{{M}^{\mathrm{op}}}m)\cdot_{{M}^{\mathrm{op}}}s=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{{M}^{\mathrm{op}}}s).$ (3)
For ((1)) we compute that
$t\cdot_{{M}^{\mathrm{op}}}(t^{\prime}\cdot_{{M}^{\mathrm{op}}}m)$ $=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{M}t^{\prime})=(m\cdot_{M}t^{\prime})\cdot_{M}t=m\cdot_{M}(t^{\prime}\cdot_{T}t)=m\cdot_{M}(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})=$
$=(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})\cdot_{{M}^{\mathrm{op}}}m.$
The case of ((2)) is symmetric. For ((3) note that
$(t\cdot_{{M}^{\mathrm{op}}}m)\cdot_{{M}^{\mathrm{op}}}s=(mt)\cdot_{{M}^{\mathrm{op}}}s=s(mt)=(sm)t=t\cdot_{{M}^{\mathrm{op}}}(sm)=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{{M}^{\mathrm{op}}}s).$
Proposition 7.
For $S$ a semiring, ${S}^{\mathrm{op}}$ has the structure of an $(S,S)$-bimodule (and hence also of either a left or a right semiring module over $S$).
Proof.
To avoid confusion, we shall denote multiplication in $S$ by $\cdot_{S}$, multiplication in ${S}^{\mathrm{op}}$ by $\cdot_{{S}^{\mathrm{op}}}$, the left action of $S$ on ${S}^{\mathrm{op}}$ by $\cdot_{L}$, and the right action of $S$ on ${S}^{\mathrm{op}}$ by $\cdot_{R}$. Given $a,b\in S$, define the left and right actions of $S$ by
$a\cdot_{L}b=a\cdot_{R}b=b\cdot_{{S}^{\mathrm{op}}}a.$
Because the two actions are given by the same operation, the fact that they commute follows immediately from associativity of multiplication in ${S}^{\mathrm{op}}$. We check that multiplication in $S$ distributes over the right action of $S$; the case of the left action is symmetric. For $a,b,c\in S$, compute
$a\cdot_{R}(b\cdot_{S}c)=(b\cdot_{S}c)\cdot_{{S}^{\mathrm{op}}}a=(c\cdot_{{S}^{\mathrm{op}}}b)\cdot_{{S}^{\mathrm{op}}}a=c\cdot_{{S}^{\mathrm{op}}}(b\cdot_{{S}^{\mathrm{op}}}a)=c\cdot_{R}(b\cdot_{R}a).$
3 Kleene Algebras and Kleene Modules
A semiring $S$ is additively idempotent if and only if for every $\in S, x+x$=xx+x=x. Additively idempotent semirings (and in fact idempotent commutative monoids more generally) [^1] Commutative idempotent monoids are easily seen to be definitionally equivalent to join-semilattices come equipped with a partial order defined by $x\leq y$ if and only if $x+y=y$. It is easy to check that $x+y$ is then the least upper bound of $x$ and $y$ in the order.
Definition 8.
A Kleene algebra $\langle K,+,\cdot,0,1,*\rangle$ is an additively idempotent semiring with an additional unary operation * that satisfies the following axioms for any $x,y,z\in K$:
  • $1+xx^{*}=1+x^{*}x=x^{*}$,
  • if $x+yz\leq z$ then $y^{*}x\leq z$,
  • if $x+zy\leq z$ then $xy^{*}\leq z$.
Note that, since semirings are defined to be unital, if a semiring $S$ is additively idempotent then any left or right module over $S$ is also additively idempotent. The argument for a left module $M$ over $S$ will be sufficient, because the case of a right module is symmetric. For $m\in M$,
$m+m=1m+1m=(1+1)m=1m=m.$
Also note that by the symmetry of the conditions on the asterate, the opposite semiring of a Kleene algebra becomes a Kleene algebra when endowed with the same asterate.
Definition 9.
Fix a semiring $S$, and let $M$ and $N$ be left $S$-modules. A map $\phi:M\rightarrow N$ is a homomorphism of left $S$-modules if and only if for every $s\in S$ and $m,n\in M$,
$\phi(m+n)=\phi(m)+\phi(n)\quad\text{and}\quad\phi(sm)=s\phi(m);$
the definitions of homomorphisms of right $S$-modules and of $(S,T)$-bimodules for $T$ a semiring are analogous. The set of all $S$-module homomorphisms from $M$ to $N$ is denoted $\operatorname{Hom}(M,N)$. Note that this set naturally comes equipped with pointwise addition; later we shall see how it can be endowed with module structure when $M$ and $N$ have additional structure. When $M=N$ we write $\mathop{\text{End}}(M)$ for $\operatorname{Hom}(M,M)$ and speak of endomorphisms of the semiring module $M$. The category of left $S$-modules and homomorphisms [^2] With functional composition as the composition operation. is denoted $\text{Mod}_{S}$, the category of right $S$-modules and homomorphisms is denoted $_{S}\text{Mod}$}, and the category of $(S,T)$-bimodules and homomorphisms is denoted $_{S}\text{Mod}_{T}$}. In all these categories isomorphism has its usual, categorical meaning (for which see e.g. [16]).
Definition 10.
Let $K$ be a Kleene algebra. A left Kleene module over $K$ is a semiring module $M$ over the semiring reduct of $K$ satisfying the following implication for $m\in M$ and $a\in K$:
$am\leq m\Rightarrow a^{*}m\leq m.$
A right Kleene module over $K$ is similarly a right semiring module $M$ satisfying $ma\leq m\Rightarrow ma^{*}\leq m$ for $m\in M$ and $a\in K$. A Kleene $(K,L)$-bimodule is a semiring $(K,L)$-bimodule which is simultaneously a left Kleene module over $K$ and a right Kleene module over $L$.
Kleene modules are a quasivariety (see e.g. [10], which gives a universal Horn axiomatization), and so by universal algebra there are free Kleene $K$-modules over arbitrary sets of generators for $K$ a fixed Kleene algebra (see for example [1]).
3.1 Examples
Algebras. Any Kleene algebra $K$ is a $(K,K)$-bimodule over itself, and a $(2,2)$-bimodule over the Kleene algebra $2$ with two elements. Moreover it is an $(A,A)$-bimodule over any subalgebra $A$ of $K$. In particular, if $K$ is a Kleene algebra with tests (as defined in [14]; a KAT), and $B$ is its distinguished Boolean subalgebra, then $B$ is a Kleene subalgebra of $K$ and $K$ is a $(B,B)$-bimodule.
Ideals. Let $I$ be a left ideal of a Kleene algebra $K$. Then $I$ is a left Kleene module with action given by multiplication. To see this, note that $I$ is clearly a left module over an additively idempotent semiring under the action of multiplication, and that for $m\in I$, if $am\leq m$ then by the *-axioms for Kleene algebras, $a^{*}m\leq m$. Similary, if $J$ is a right ideal of $K$ then $J$ is a right Kleene module, and if $J$ is a two-sided ideal then it is a Kleene $(K,K)$-bimodule.
Quotients. Let $\equiv$ be a congruence of a Kleene algebra $K$, and set $Q=K/{\equiv}$. If $Q$ is in fact a Kleene algebra, then $Q$ is a Kleene $K$-bimodule with action given by multiplication. This follows from the fact that $\equiv$ is a congruence, and provides an example of a Kleene algebra construction that comes with a natural Kleene bimodule structure.
Function modules. Let $K$ be a Kleene algebra, $B$ be a set, and consider the set $K^{B}$ of functions from $B$ to $K$ endowed with pointwise addition and zero as the constant-zero function. Define left and right actions of $K$ on $K^{B}$ by $(am)(b)=am(b)$ and $(ma)(b)=m(b)a$ for $a\in K$, $m\in K^{B}$. It is immediate that these operations give a $(K,K)$-bimodule structure on $K^{B}$. Moreover, if $B$ is finite, then $K^{B}$ has the usual finite basis given by characteristic functions for singletons in $B$, and hence is the free $(K,K)$-bimodule with generators $B$. Forgetting either the right or the left action, it is also the free left and the free right $K$-module with generators $B$. One can also restrict from $K^{B}$ to functions supported on a set in some lattice-ideal for $\operatorname{\mathcal{P}}(B)$. It is easy to see that the finitely-supported [^3] So, supported in the ideal of finite sets. elements of $K^{B}$ form the free $K$-module with generators $B$.
Homomorphism and endomorphsim modules. Fix Kleene algebras $K$ and $S$ and let $M$ and $N$ be $(K,S)$-bimodules. Denote by $H$ the set $\operatorname{Hom}(M,N)$, which we shall soon endow with additional structure. For $\alpha,\beta\in\operatorname{Hom}(M,N)$, define $\alpha+\beta$ to be the pointwise sum of $\alpha$ and $\beta$ and $0_{H}$ to be the constant function with value zero. This trivially gives $\operatorname{Hom}(M,N)$ the structure of a commutative monoid. Now define a left action of $S$ by $(s\cdot_{H}\alpha)(m)=\alpha(m\cdot_{M}s)$ for $s\in S$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$; it is straightforward to check that this gives $\operatorname{Hom}(M,N)$ the structure of a left module over $S$. It remains to check for $s\in S$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$ that if $s\alpha\leq\alpha$ then $s^{*}\alpha\leq\alpha$. Assume that $s\alpha\leq\alpha$, so for each $m\in M$, $s\alpha(m)\leq\alpha(m)$. Applying the action gives that $\alpha(m\cdot_{M}s)=\alpha(m)\cdot_{N}s\leq\alpha(m)$, so because $M$ is a right semiring module over $S$, $\alpha(m)\cdot_{N}s^{*}=\alpha(m\cdot_{M}s^{*})\leq\alpha(m)$, which means $s^{*}\cdot_{H}\alpha(m)\leq\alpha(m)$, that is, $s^{*}\cdot_{H}\alpha\leq\alpha$.
Reversing order when appropriate handles the case of homomorphisms of into right modules Specifically, if $M$ and $N$ are $(S,K)$-bimodules, then $\operatorname{Hom}(M,N)$ becomes a right $K$-module when endowed with pointwise addition, the constant function zero, and the right action of $K$ defined by $(\alpha\cdot_{H}k)(m)=\alpha(k\cdot_{M}m)$ for $k\in K$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$.
Let $K$ be a Kleene algebra, $M$ a $(K,K)$-bimodule. Define
$\mathop{\text{End}}(M)=\operatorname{Hom}(M,M),$
the endomorphisms of $M$. By the preceding example $\mathop{\text{End}}(M)$ is a $(K,K)$-bimodule over the semiring reduct of $K$. Now we verify that it in fact carries the structure of a Kleene $(K,K)$-bimodule. For $k\in K$ and $\alpha\in\mathop{\text{End}}(M)$, if $\alpha k\leq\alpha$, then for every $m\in M$, $\alpha(m)k\leq\alpha(m)$, so $\alpha(m)k^{*}\leq\alpha(m)$ for every $m\in M$ and hence $\alpha k^{*}\leq\alpha$. Consequently $\mathop{\text{End}}(K)$ is a right Kleene module over $K$. A similar argument gives that $\mathop{\text{End}}(K)$ is a left Kleene module over $K$, so it remains to verify that the left and right actions of $K$ commute. This is given by the following computation, where $s,k\in K$, $m\in M$, and $\alpha\in\mathop{\text{End}}(K)$.
$(s\alpha)k](m)=(s\alpha)(km)=\alpha(kms)=(\alpha k)(ms)=[s(\alpha k)](m).$
In the special case of $K^{n}$ viewed as a $({K}^{\mathrm{op}},{K}^{\mathrm{op}})$-bimodule with pointwise scalar multiplication, ${\mathop{\text{End}}(M)}^{\mathrm{op}}$ is the $(K,K)$-bimodule of $n\times n$ matrices over $K$. The reason for taking the opposite is that the left and right actions of $K$ on $M$ are reversed in the definition of $\operatorname{Hom}(M,M)=\mathop{\text{End}}(K)$.
Also, as in the classical theory of modules over rings, the dual $M^{\circ}$ of a Kleene $(K,K)$-bimodule $M$ is the Kleene $(K,K)$-bimodule $\operatorname{Hom}(M,{{}_{K}}K_{K})$. Since the concepts of homomorphism module, endomorphism module, and the dual of a module will be used below, we recapitulate the definitions here.
Definition 11.
Let $K$, $S$ be Kleene algebras and $M$, NN Klene $(K,S)$-bimodules. The homomorphsim module is the set of homomorphisms from $M$ to $N$, endowed with the operations defined at the beginning of this example. If $K=S$, the endomorphism module of $M$ is $\mathop{\text{End}}(M)=\operatorname{Hom}(M,M)$, and the dual of $M$ is $M^{\circ}=\operatorname{Hom}(M,{{}_{K}}K_{K})$.
Propositional Dynamic Logic. Fix sets $P$ of action atoms and $T$ of test atoms, respectively. [^4] For details concerning propositional dynamic logic see for instance [13]. Let $K$ be the free Kleene algebra generated by $P$, and $B$ the free Boolean algebra generated by $T$. The Lindenbaum-Tarski algebra $A$ of propositional dynamic logic with action atoms $P$ and test atoms $T$ can be viewed as a $(B,B)$-bimodule via the subalgebra bimodule structure, but also has the structure of a left $K$-module via the action $pa=[\langle p\rangle\phi]$ where $p\in K$, $a\in A$, $a=[\phi]$, and brackets denote the equivalence class of a formula in the Lindenbaum-Tarski algebra. This follows immediately from the axioms for propositional dynamic logic.
3.2 Tensor products of Kleene modules
Tensor products of modules are defined so that the tensor product functor $\otimes N$ is left-adjoint to the hom-functor $\operatorname{Hom}(N,{-})$ for each module $N$ of the correct type. This is constructed by taking the quotient of formal sums of pairs from ${{}_{A}}M_{B}\times{{}_{B}}N_{C}$ by the least congruence imposing $A$-linearity in the first coordinate, $C$-linearity in the second coordinate, and balance, which means that for any $m\in M$, $n\in N$, and $b\in B$, $(mb,n)$ is identified with $(m,bn)$. But in the context of Morita equivalence only the isomorphism class of tensor products is relevant, so we define what it means to be a tensor product via a categorical adjunction requirement and then check that the definition just outlined satisfies this requirement.
Definition 12.
Fix Kleene algebras $A$, $B$, $C$. A tensor product is a functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$ such that for each $N\in{{}_{B}\text{Mod}_{C}}$, the functors $M.\mapsto F(M,N)$ and $P\mapsto\operatorname{Hom}(N,P)$ form the left and right parts of an adjunction, respectively.
Note that it follows from the universal property for adjoints that tensor products, if they exist, are unique up to isomorphism [16, IV.1]. One can also define tensor products producing left and right Kleene modules, with types ${{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}}\rightarrow{{}_{A}\text{Mod}}$ and $\text{Mod}_{A}\times{{}_{A}\text{Mod}_{B}}\rightarrow\text{Mod}_{A}$, in the obvious way. We omit the definition since it will not be used in the sequel.
Defining tensor products of Kleene modules via an adjunction requirement does not guarantee they exist; this is established next.
Proposition 13.
For any Kleene algebras $A$, $B$, $C$ there is a tensor product functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$.
Proof.
We proceed by direct construction. Fix $M\in{{}_{A}\text{Mod}_{B}}$ and $N\in{{}_{B}\text{Mod}_{C}}$. Let $\Xi$ be the Kleene $(A,C)$-bimodule freely generated by $M\times N$, and let $\Phi$ be the congruence generated by the relations
$(m+m^{\prime},n)\mathrel{\Phi}(m,n)+(m^{\prime},n),(m,n+n^{\prime})\mathrel{\Phi}(m,n)+(m,n^{\prime}),(mb,n)\mathrel{\Phi}(m,bn)$
for $m,m^{\prime}\in M$, $n,n^{\prime}\in N$, and $b\in B$. Define $M\otimes_{B}N=\Xi/\Phi$. This immediately has the structure of a semiring because semirings are varieties and so any quotient by a semiring congruence is a semiring (see, for instance, [3]). To see that this quotient in fact defines a Kleene module, we must check for $a\in A$, $c\in C$, $m\in M$, and $n\in N$ that if $a(m,n)\leq(m,n)$ then $a^{*}(m,n)\leq(m,n)$, and that if $(m,n)c\leq(m,n)$ then $(m,n)c^{*}\leq(m,n)$. These inequalities follow immediately from the fact that $M$ is a left module over $A$ and $N$ is a right module over $M$, together with the fact that $(m,n)+(m,n)=(m,n)$.
It suffices to establish that the functors $M\mapsto M\otimes_{B}N$ and $P\mapsto\,\operatorname{Hom}(N,P)$ are adjoint. This means that there is a natural isomorphism $\operatorname{Hom}(M\otimes_{B}N,P)\cong\operatorname{Hom}(M,\operatorname{Hom}(N,P))$ for $M\in{{}_{A}\text{Mod}_{B}}$ and $N,P\in{{}_{B}\text{Mod}_{C}}$. For $\varphi\in\operatorname{Hom}(M\otimes_{B}N,P)$ define $\overline{\varphi}$ by $\overline{\varphi}(m)(n)=\varphi(m\otimes n)$. This defines a homomorphism $M\rightarrow\operatorname{Hom}(N,P)$ by freeness of $\Xi$ and the construction of the congruence $\Phi$. Let $M^{\prime}\in{{}_{A}\text{Mod}_{B}}$, $P^{\prime}\in{{}_{B}\text{Mod}_{C}}$, $\alpha\mathrel{:}M^{\prime}\rightarrow M$, and $\beta\mathrel{:}P\rightarrow P^{\prime}$. To see that the map $\varphi\mapsto\overline{\varphi}$ is a natural transformation, we must verify that the following diagram commutes, where the horizontal arrows represent application of the transformation $\varphi\mapsto\overline{\varphi}$.
$\operatorname{Hom}(M\otimes_{B}N,P)$$\operatorname{Hom}(M,\operatorname{Hom}(N,P))$$\operatorname{Hom}(M^{\prime}\otimes_{B}N,P^{\prime})$$\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$$\scriptstyle{\operatorname{Hom}(\alpha\otimes\text{id},\beta)}$$\scriptstyle{\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))}$
Let $\varphi\in\operatorname{Hom}(M\otimes_{B}N,P)$. Then for $m^{\prime}\in M^{\prime}$ and $n\in N$,
$\overline{\operatorname{Hom}(\alpha\otimes\text{id},\beta)(\varphi)}(m^{\prime})(n)$ $=(\beta\circ\varphi\circ(\alpha\otimes\text{id}))(m^{\prime}\otimes n)$
$=\beta(\varphi((\alpha\otimes\text{id})(m^{\prime}\otimes n)))$
$=\beta(\varphi(\alpha(m^{\prime})\otimes n))$
and
$\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))(\overline{\varphi})(m^{\prime})(n)$ $=\operatorname{Hom}(N,\beta)(\overline{\varphi}(\alpha(m^{\prime}))(n))$
$=\operatorname{Hom}(N,\beta)(\varphi(\alpha(m^{\prime})\otimes n))$
$=\beta(\varphi(\alpha(m^{\prime})\otimes n)).$
Consequently $\varphi\mapsto\overline{\varphi}$ is a natural transformation.
It remains to find an inverse to the natural transformation $\varphi\mapsto\overline{\varphi}$. The obvious candidate is given for $M\in{{}_{A}\text{Mod}_{B}}$, $P\in{{}_{B}\text{Mod}_{C}}$, $\psi\in\operatorname{Hom}(M,\operatorname{Hom}(N,P))$, $m\in M$, and $n\in N$ by $\hat{\psi}(m\otimes n)=\psi(m)(n)$. It is immediate that the transformation $\psi\mapsto\hat{\psi}$ is inverse to the transformation $\phi\mapsto\overline{\phi}$. It remains to verify that $\psi\mapsto\hat{\psi}$ is a natural transformation. For this we need to check that the following diagram commutes, where $M,M^{\prime}\in{{}_{A}\text{Mod}_{B}}$, $P,P^{\prime}\in{{}_{B}\text{Mod}_{C}}$, $\alpha\mathrel{:}M\rightarrow M^{\prime}$, $\beta\mathrel{:}P^{\prime}\rightarrow P$, and the horizontal arrows are given by application of the transformation $\psi\mapsto\hat{\psi}$.
$\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$$\operatorname{Hom}(M^{\prime}\otimes_{B}N,P^{\prime})$$\operatorname{Hom}(M,\operatorname{Hom}(N,P))$$\operatorname{Hom}(M\otimes_{B}N,P)$$\scriptstyle{\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))}$$\scriptstyle{\operatorname{Hom}(\alpha\otimes\text{id},\beta)}$
Let $\psi\in\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$. Then for $m\otimes n\in M\otimes_{B}N$,
[…]
$=(\operatorname{Hom}(N,\beta)\circ\psi\circ\alpha)(m)(n)$
$=\operatorname{Hom}(N,\beta)(\psi(\alpha(m)))(n)$
$=(\beta\circ\psi(\alpha(m)))(n)$
$=\beta(\psi(\alpha(m))(n))$
One more computation now verifies that this diagram commutes:
$\operatorname{Hom}(\alpha\otimes\text{id},\beta)(\hat{\psi})(m\otimes n)$ $=(\beta\circ\hat{\psi}\circ(\alpha\otimes\text{id}))(m\otimes n)$
$=\beta(\hat{\psi}(\alpha(m)\otimes n))$
$=\beta(\psi(\alpha(m))(n)).$
Therefore the transformation $\psi\mapsto\hat{\psi}$ is inverse to the natural transformation $\varphi\mapsto\overline{\varphi}$, so these transformations are natural isomorphisms. ∎
The next results will be needed to verify that the Morita skeleton category from which we define Morita equivalence is indeed a category.
Proposition 14.
Let $A$, $B$, $C$, $D$ be Kleene algebras, and ${{}_{A}}M_{B}$, ${{}_{B}}N_{C}$, ${{}_{C}}P_{D}$ be Kleene bimodules. Then
$({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D}\cong{{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D}),$
and also ${{}_{A}}A_{A}\otimes{{}_{A}}M_{B}\cong{{}_{A}}M_{B}$ and ${{}_{A}}M_{B}\otimes{{}_{B}}B_{B}\cong{{}_{A}}M_{B}$.
Proof.
For the isomorphism $({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D}\cong{{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D})$ it suffices to check that the correspondence
$(m\otimes n)\otimes p\mapsto m\otimes(n\otimes p)$
for $m\in{{}_{A}}M_{B}$, $n\in{{}_{B}}N_{C}$, and $p\in{{}_{C}}P_{D}$; is well-defined, for then it extends linearly to a homomorphism which clearly has as inverse the homomorphism obtained from the (similarly well-defined) correspondence
$m\otimes(n\otimes p)\mapsto(m\otimes n)\otimes p.$
Suppose then that $m,m^{\prime}\in M$, $n,n^{\prime}\in N$, and $p,p^{\prime}\in P$, and that $(m\otimes n)\otimes p=(m^{\prime}\otimes n^{\prime})\otimes p^{\prime}$. This is witnessed by a sequence of steps of applying additivity in each variable and commuting scalars across the tensor product, that is to say, applying the generators of the congruence used to define the tensor product. Because scalars multiplying the middle variable may be pulled out of the inner tensor product (that is, $m^{\prime\prime}\otimes n^{\prime\prime}c=(m^{\prime\prime}\otimes n^{\prime\prime})c$ and $bn^{\prime\prime}\otimes p^{\prime\prime}=b(n^{\prime\prime}\otimes p^{\prime\prime})$ and then moved across the outer tensor product, the same sequence of operations can be applied to $m\otimes(n\otimes p)$ to obtain $m^{\prime}\otimes(n^{\prime}\otimes p^{\prime})$. This shows that the correspondence is well-defined and completes the proof of associativity of the tensor product up to isomorphism. The isomorphsims ${{}_{A}}A_{A}\otimes{{}_{A}}M_{B}\cong{{}_{A}}M_{B}$ and ${{}_{A}}M_{B}\otimes{{}_{B}}B_{B}\cong{{}_{A}}M_{B}$ are obtained by linearly extending the maps $1_{A}\otimes m\mapsto m$ and $m\otimes 1_{B}\mapsto m$ to homomorphisms and noting that these have respective inverses determined by $m\mapsto 1_{A}\otimes m$ and $m\mapsto 1_{B}\otimes B$. ∎
4 The Morita Skeleton of Semirings and of Kleene Algebras
The case of Morita equivalence for Kleene algebras is a straightforward adaptation of that for rings.
Definition 15.
The Morita skeleton of semirings is a $1$-category with objects isomorphism classes of semirings, [^5] Scott’s trick [17] may be used to represent these equivalence classes as sets. and for $A$ and $B$ semiringss the morphisms between them are all isomorphism classes of semiring $(A,B)$-bimodules (so this category is not locally small). The identity at an algebra $A$ is ${{}_{A}}A_{A}$, and composition of morphisms is given by the operation on isomorphism classes induced by tensor product (which is clearly well-defined because the tensor product is defined via an adjunction). Semirings $A$ and $B$ are called Morita-equivalent precisely when they are isomorphic in the Morita skeleton of semirings.
The Morita skeleton of Kleene algebras is the subcategory of the Morita skeleton of semirings with objects isomorphisms of Kleene algebras and morphisms isomorphism classes of Kleene bimodules.
In the case of rings one easily obtains a bicategory structure by taking $2$-cells to be homomorphisms between modules [2], and presumably this can be generalized to semirings and adapted to Kleene algebras, but since we do not use this extra structure we do not develop it.
The Morita skeleton (of semirings or of Kleene algebras) extends the usual category of semirings (Kleene algebras) and homomorphisms via the following notion of homomorphism module, which realizes homomorphisms as particular bimodules.
Definition 16.
Let $A$, $B$ be semirings (Kleene algebras) and $h\mathrel{:}A\rightarrow B$ be a homomorphism. The homomorphism module $E_{h}$ is a semiring (Kleene) $(A,B)$-bimodule whose elements and addition are given by the additive reduct of $B$, with the following multiplication actions on $b\in B$: For $a\in A$, $a\cdot_{E_{h}}b=h(a)\cdot_{B}b$. For $b^{\prime}\in B$, $b\cdot_{E_{h}}b^{\prime}=b\cdot_{B}b^{\prime}$.
The next proposition verifies that $E_{h}$ has the structure of a semiring (Kleene) bimodule.
Proposition 17.
Let $A$ and $B$ be semirings (Kleene algebras) and $h:A\rightarrow B$ be a homomorphism. Then $E_{h}$ satisfies the axioms of a semiring (Kleene) $(A,B)$-bimodule.
Proof.
Clearly $E_{h}$ has the structure of a commutative idempotent monoid. The axioms for the right action of $B$ are immediate, and the axioms for the left action of $A$ follow because $h$ is a homomorphism. The fact that the two actions commute is an instance of associativity in $B$. In case $A$ and $B$ are Kleene algebras, the asterate inequality for the right action follows from the fact that $B$ is a Kleene algebra, while the asterate inequality on the left is verified by the following computation, where $a\in A$ and $b\in B$. If $a\cdot_{E_{h}}b=h(a)b\leq b$, then $h(a)^{*}b\leq b$ because $B$ is a Kleene algebra, and hence $a^{*}\cdot_{E_{h}}b=h(a^{*})b\leq b$ because $h(a)^{*}=h(a^{*})$ since $h$ is a homomorphism of Kleene algebras. ∎
Next we verify that the tensor product of homomorphism bimodules corrsponds in the Morita skeleton (of semirings or of Kleene algebras) to the homomorphism module of the composition, so that homomorphism modules do indeed give an extension of the category of semirings (Kleene algebras) and homomorphisms to the Morita skeleton.
Proposition 18.
Let $A$, $B$, $C$ be semirings (Kleene algebras) and $f\mathrel{:}A\rightarrow B$ and $g\mathrel{:}B\rightarrow C$ be homomorphisms. Then $E_{g\circ f}\cong E_{f}\otimes E_{g}$.
Proof.
We shall show that the map $\phi\mathrel{:}E_{g\circ f}\rightarrow E_{f}\otimes E_{g}$ given by $\phi(c)=1\otimes c$ is an isomorphism. To see that this is a homomorphism, additivity and the commutativity with the right action of $C$ are immediate. For the left action of $A$, note that for $a\in A$,
$\phi(a\cdot_{E_{g\circ f}}c)$ $=1\otimes(g\circ f)(a)\cdot_{C}c$
$=f(a)\otimes c$
$=a(1\otimes c)=a\phi(c).$
Now consider $b\otimes c\in E_{g}\otimes E_{f}$. Define $\psi\mathrel{:}E_{f}\otimes E_{g}\rightarrow E_{g\circ f}$ by $\psi(b\otimes c)=b\cdot{E_{g}}c$, and note that this is well-defined because $b\otimes c=1\otimes b\cdot_{E_{g}}c$ for $b\in B$, $c\in C$. It is clearly additive and commutes with the right action of $C$, so we demonstrate that it commutes with the left action of $A$. Let $a\in A$ and note that
$\psi(a(b\otimes c))$ $=\psi(f(a)(b\otimes c))$
$=\psi(f(a)\otimes b\cdot_{E_{g}}c)$
$=f(a)(b\cdot_{E_{g}}c)$
$=a\psi(b\otimes c).$
It remains to compute
$\phi(\psi(b\otimes c))=\phi(b\cdot_{E_{g}}c)=1\otimes(b\cdot_{E_{g}}c)=b\otimes c$
and
$\psi(\phi(c))=\psi(1\otimes c)=1\cdot_{E_{g}}c=c$
to see that $\phi$ and $\psi$ are inverse bimodule homomorphisms, hence isomorphisms. Therefore $E_{g\circ f}\cong E_{f}\otimes E_{g}$. ∎
Another way of expressing Morita equivalence for rings is as equivalence of the corresponding module categories. This carries over to Kleene algebras, as we now argue. If $K$ is Morita-equivalent with $S$ via bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$, then for any left $K$-module ${{}_{K}}P$,
$M\otimes(N\otimes P)\cong(M\otimes N)\otimes P\cong K\otimes P\cong P,$
and for any left $S$-module ${{}_{S}}Q$,
$N\otimes(M\otimes Q)\cong(N\otimes M)\otimes Q\cong S\otimes Q\cong Q,$
so the functors $P\mapsto N\otimes P$ from $\text{Mod}_{K}$ to $\text{Mod}_{S}$, and $Q\mapsto N\otimes Q$ from $\text{Mod}_{S}$ to $\text{Mod}_{K}$, constitute an equivalence of categories. The right-module categories $_{K}\text{Mod}$} and $_{S}\text{Mod}$} are equivalent by an exactly symmetric argument, and combining these arguments together gives that the bimodule categories $_{K}\text{Mod}_{S}$} and $_{S}\text{Mod}_{K}$} are equivalent.
As in the classical case, a Kleene algebra is Morita-equivalent with each of its finite-dimensional matrix rings. Because deterministic finite automata relative to a Kleene algebra $K$ correspond to square matrices over $K$, together with vectors of start and end states [11, Definition 12], this shows that reducts of automata in one Kleene algebra (forgetting the start and end states) correspond to single elements in a Morita-equivalent Kleene algebra. Unfortunately the author has not yet satisfactorily investigated concrete results which can be proved using this correspondence.
Proposition 19.
Let $K$ be a Kleene algebra, $n$ a positive integer, and $M$ be the Kleene algebra of $n\times n$ matrices over $K$. Then $K$ is Morita-equivalent with $M$.
Proof.
The $(M,K)$-bimodule $K^{n}$ and its dual $K^{n\circ}$ will witness this equivalence. First consider $K^{n}\otimes K^{n\circ}$. For $1\leq i\leq n$ let $e_{i}$ be the “$i$th unit vector” of $K^{n}$, that is the tuple which is zero everywhere except position $i$, where it is $1$. Note that the set $\{e_{i}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n}$. Let $e_{i}^{\circ}$ be the dual of this module element, that is, $e_{i}^{\circ}(e_{i})=1$ and $e_{i}^{\circ}(e_{j})=0$ for $j\neq i$. It is easy to show that the set $\{e_{i}^{\circ}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n\circ}$. For $1\leq i,j\leq n$, let $E_{ij}\in M$ be the matrix which is zero everywhere except at position $(i,j)$, where it is $1$. Note that these freely generate $M$, and that the set $\{e_{i}\otimes e_{j}^{\circ}\mathrel{:}1\leq i,j\leq n\}$ freely generates $K^{n}\otimes K^{n\circ}$, because only elements of $K$ can be transported across the tensor product. Define a map $\phi\mathrel{:}K^{n}\otimes K^{n\circ}\rightarrow M$ by $\phi(e_{i}\otimes e_{j}^{\circ})=E_{ij}$; we show that $\phi$ is an isomorphism. Its inverse, naturally, will be given by $\psi(E_{ij})=e_{i}\otimes e_{j}^{\circ}$. By free generation $\phi$ and $\psi$ are well-defined homomorphisms, and it is immediate that they are inverses. Consequently $K^{n}\otimes K^{n\circ}\cong M$, which is half of a Morita equivalence.
For the other half we look at $K^{n\circ}\otimes K^{n}$, and our goal is to show this is isomorphic to $K$. For $e_{i}^{\circ}\otimes e_{j}\in K^{n\circ}\otimes K^{n}$, define $\alpha(e_{i}^{\circ}\otimes e_{j})$ to give the $ij$-entry of the identity matrix over $K$; that is, 0 if $i\neq j$ and $1$ if $i=j$. By free generation this determines a homomorphism $\alpha\mathrel{:}K^{n\circ}\otimes K^{n}\rightarrow K$. Define also $\beta\mathrel{:}K\rightarrow K^{n\circ}\otimes K^{n}$ by $\beta(a)=a(e_{1}^{\circ}\otimes e_{1})$ for $a\in A$. To complete the proof we show that $\alpha$ and $\beta$ are inverses. Certainly for $a\in A$ we have that
$\alpha(\beta(a))=\alpha(a(e_{1}^{\circ}\otimes e_{1}))=a\alpha(e_{1}^{\circ}\otimes e_{1})=a1=a.$
To see that $\beta\circ\alpha=\text{id}$ it suffices to show that for any $i\leq n$, $e_{i}^{\circ}\otimes e_{i}=e_{1}^{\circ}\otimes e_{1}$ and for $j\leq n$, $j\neq i$, $e_{i}^{\circ}\otimes e_{j}=0$. For the latter let $E_{j}$ be the matrix obtained by replacing the $j$th row of the zero matrix by the “row vector” $e_{j}^{\circ}$. One computes easily that $E_{j}e_{j}=e_{j}$, and that $e_{i}^{\circ}E_{j}=0$. Hence
$e_{i}^{\circ}\otimes e_{j}=e_{i}^{\circ}\otimes E_{j}e_{j}=e_{i}^{\circ}E_{j}\otimes e_{j}=0\otimes e_{j}=0.$
For the case of $i=j$ we show that
$e_{i}^{\circ}\otimes e_{i}=\left(\sum_{i=1}^{n}e_{i}^{\circ}\right)\otimes\left(\sum_{i=1}^{n}e_{i}\right).$
Before writing this out in generality, an example might be instructive. We compute in the case $n=3$ as follows:
$\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$ $=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}=\begin{pmatrix}1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
To compute this in generality, let $\overline{1}$ be the $n\times n$ matrix with every entry $1$, and $\overline{e}_{i}$ be the result of replacing the $i$th row of $\overline{1}$ with the entries of $e_{i}$. It is straightforward to compute the following:
  • $e_{i}^{\circ}\overline{e}_{i}=e_{i}^{\circ}$.
  • $\overline{e}_{i}e_{i}=\sum_{i=1}^{n}e_{i}$.
  • $\overline{1}\sum_{i=1}^{n}e_{i}=\sum_{i=1}^{n}e_{i}$.
  • $e_{i}^{\circ}\overline{1}=\sum_{i=1}^{n}e_{i}^{\circ}$.
Then we have
$e_{i}^{\circ}\otimes e_{i}$ $=e_{i}^{\circ}\overline{e}_{i}\otimes e_{i}=e_{i}^{\circ}\otimes\overline{e}_{i}e_{i}=e_{i}^{\circ}\otimes\sum_{i=1}^{n}e_{i}$
$=e_{i}^{\circ}\otimes\overline{1}\sum_{i=1}^{n}e_{i}=e_{i}^{\circ}\overline{1}\otimes\sum_{i=1}^{n}e_{i}=\left(\sum_{i=1}^{n}e_{i}^{\circ}\right)\otimes\sum_{i=1}^{n}e_{i}.$
This completes the proof. ∎
There is a well-known characterization of Morita equivalence for rings which extends to semirings [15, 9]. To state it we need the following definition:
Definition 20.
Let $R$ be a semiring and $e\in R$. The element $e$ is multiplicatively idempotent precisely when $e^{2}=e$, and full if and only if $ReR=R$.
Proposition 21 ([9]).
Semirings $R$ and $S$ are Morita-equivalent if and only if one of the following occurs for some dimension $n$:
  1. 1.
    for some full idempotent $e$ in the semiring $M$ of $n\times n$ matrices over the semiring $R$, $S\cong eMe$, or else
  2. 2.
    for some full idempotent $e^{\prime}$ in the semiring $M^{\prime}$ of $n\times n$ matrices over the semiring $S$, $R\cong e^{\prime}M^{\prime}e^{\prime}$.
Certainly Morita-equivalent Kleene algebras are Morita equivalent as semirings, and so any Kleene algebra $S$ Morita-equivalent to the algebra $K$ is, up to isomorphism, of the form $eMe$ for some finite-dimensional matrix ring $R$ over $K$ and some full idempotent $e\in M$. It is natural to ask whether Kleene algebras whose semiring reducts are Morita equivalent must be Morita equivalent as Kleene algebras. That is to say, if there is a semimodule Morita isomorphism between Kleene algebras $K$ and $S$, must there also be a Kleene algebra Morita isomorphism?
Theorem 22 (Morita rigidity of Kleene algebras over semirings).
Let $K$ and $S$ be a Kleene algebras. If the semiring reducts of $K$ and $S$ are Morita-equivalent via semiring bimodules, then in fact $K$ and $S$ are Morita-equivalent via Kleene bimodules..
The statement of this theorem is exactly what we mean when stating that the category of Kleene algebras is Morita-rigid with respect to the category of semirings.
Proof.
Using proposition 21, we may assume without loss of generality that for some dimension $n$, $S$ is isomorphic to $eMe$ for some full idempotent $e$ of the Kleene algebra $M$ of matrices over $S$. The asterate of $M$ is the standard matrix asterate for Kleene algebras; see [11] for the definition and basic properties. Note that $K^{n\circ}e\otimes eK^{n}$ is a submodule of $K^{n\circ}\otimes K^{n}$, so because the idempotent $e$ is full, in fact $K^{n\circ}e\otimes eK^{n}\cong K$. Conversely,
$eK^{n}\otimes K^{n\circ}e\cong e(K^{n}\otimes K^{n\circ})e\cong eMe\cong S.$
Therefore the Kleene modules $K^{n\circ}e$ and $eK^{n}$ witness that $K$ and $S$ are Morita-equivalent. ∎
5 Conclusion and further directions
At the time of this writing Morita equivalence is a new theoretical tool being introduced into the study of Kleene algebras and related structures, but it does have some intriguing potential for applications to the theory of programming language semantics. One observation is that it is possible to freely change scalars from one Kleene-type algebra $K$ to a Morita equivalent one $S$, because if bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$ witness the Morita equivalence of $K$ and $S$, then for any left module ${{}_{K}}P$ we can reversibly pass to the $S$-module ${{}_{S}}P={{}_{S}}N_{K}\otimes{{}_{K}}P$, the reversibility being witnessed by the fact that
${{}_{K}}M_{S}\otimes{{}_{S}}P\cong({{}_{K}}M_{S}\otimes{{}_{S}}N_{K})\otimes{{}_{K}}P\cong{{}_{K}}K_{K}\otimes{{}_{K}}P\cong{{}_{K}}P.$
Changes of scalars for left modules and bimodules are handled similarly.
It is also of interest to investigate what additional structure can be added to a Kleene algebra (such as order-meets, to form a Kleene lattice, or order-residuals, to form an action algebra; see [12]) while preserving Morita rigidity in the sense that Morita equivalence of semiring reducts implies Morita equivalence of the full algebras (via bimodules which in some appropriate sense preserve the additional structure). This question was suggested to the author by Jose Meseguer at Logic and Applications 2025 in Dubrovník, Croatia The author has investigated this problem but removed results stated in an earlier version due to subtleties concerning homomorphism modules and tensor products pointed out by a diligent, anonymous referee.
This work is just the beginning of what could be a large research program. One exciting direction is to develop homology and cohomology theory for Kleene algebras and investigate what implications the presence of nontrivial (co)homology classes have for computations represented in a given algebra.